OpenStudy (anikate):

Write a linear function that relates the height of a bunny hop to the weight of the bike? x: 19, 21, 23 y: 10.2, 9.8, 9.4

OpenStudy (anikate):

x is weight and y is height

OpenStudy (anonymous):

Slope of the line = rise/run. (10.2-9.8)/(19-21) = 0.4/-2 = -0.2 slope. Plug in -0.2 into y = mx + b gives you y = -0.2x + b. Plug in any point (let's say 19, 10.2) and it gives you 10.2 = -0.2(19) + b. b = 14, so the answer is y = -0.2x + 14

OpenStudy (anikate):

sry how did u find B

OpenStudy (anonymous):

10.2 = -0.2(19) + b. 10.2 = -3.8 + b 14 = b

OpenStudy (anikate):

oh ok thx